OpenStudy (anonymous):

Simplify. 3 times the 3rd root of 16 5 times the 3rd root of 2 3 times the 3rd root of 2 12 times the 3rd root of 2 27 times the 3rd root of 2 Question 3 (Multiple Choice Worth 2 points) The girls group rented 2 canoes and 3 inner tubes for a total cost of $80. The boys group rented 1 canoe and 7 inner tubes for a total cost of $95. What is the cost to rent 1 canoe? $10 $15 $25 $18 Question 4 (Multiple Choice Worth 2 points) Simplify 125 to the power of negative 2 over 3 –25 25 one over 25 negative one over 25

OpenStudy (anonymous):

you should really try to do homework on your own

OpenStudy (anonymous):

whats you guesses ?

OpenStudy (anonymous):

1) b 2)c 3)a 4)a

OpenStudy (anonymous):

i think 4s wrong tho

OpenStudy (anonymous):

i think there is a mistake in question 4

OpenStudy (anonymous):

are the other 3 correct?

OpenStudy (anonymous):

-1/25

OpenStudy (anonymous):

\[\frac{125^{-2}}{3} = \frac{1}{125^2*3}\]

OpenStudy (anonymous):

then what? lol

OpenStudy (anonymous):

i think there is a mistake in the actual answer choices because it makes no since when you have a big number to the neg power your gonna get a small number..

OpenStudy (anonymous):

oh :/ these are the options -25 25 1/25 -1/25

OpenStudy (anonymous):

mathstudent will help because im confused on Q4

OpenStudy (mathstudent55):

Is the question in 4. \( \large {125^{ -\frac{2}{3} } }\) ?

OpenStudy (anonymous):

were my other three correct?

OpenStudy (anonymous):

yes :) thats the question

OpenStudy (anonymous):

oh :o

OpenStudy (mathstudent55):

\(\large {125^{ -\frac{2}{3} } }\) \( = \large \dfrac{1}{125^{ \frac{2}{3}} } \) \(= \large \dfrac{1}{ (\sqrt[3]{125})^{2} } \) \(= \large \dfrac{1}{ 5^{2} } \) \(= \large \dfrac{1}{ 25 } \)

OpenStudy (anonymous):

u mean c?

OpenStudy (mathstudent55):

Correct. For 4, the answer is C.

OpenStudy (mathstudent55):

Are you sure your answer choices for 1 are correct? I don't see the correct answer in any of the 4 choices.

OpenStudy (anonymous):

b is 6 ^3sqrt2

OpenStudy (mathstudent55):

This is problem 1, right? \( \large{3 \sqrt[3]{16} }\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its 6 ^3sqrt2 correct?

OpenStudy (mathstudent55):

Choice B above reads "3 times the 3rd root of 2"

OpenStudy (anonymous):

i know, thats wrong its supposed to be 6\[6 ^3\sqrt{2}\]

OpenStudy (mathstudent55):

\( \large{3 \sqrt[3]{16} } \) \(= \large{3 \sqrt[3]{8 \times 2} } \) \(= \large{3 \sqrt[3]{2^3 \times 2} } \) \(= \large{3 \times 2 \sqrt[3]{2} } \) \(= \large{6 \sqrt[3]{2} } \) Then it's correct. Problem 1: Answer is B: \(\large{6 \sqrt[3]{2} } \)

OpenStudy (anonymous):

yay

OpenStudy (mathstudent55):

Why did you write 4 answers above when there are only 3 questions?

OpenStudy (anonymous):

Which expression represents 8 x to the three sevenths power in radical form? 8 times the 3rd root of x to the seventh 8 times the seventh root of x to the third power the seventh root of 8 x to the third power the third root of 8 x to the seventh power I GOT A What is the equation of the line that passes through the point (4, 1) and is perpendicular to the line y = –2x + 4? y = negative one half x – 3 y = negative one half x + 1 y = one half x + 4 y = one half x – 1 EITHER A OR B Solve 6the square root of x plus one) – 3 = 15 x = 8 x = 71 x = 75 x = 107 I GOT B Which system of inequalities is graphed below?

OpenStudy (mathstudent55):

The canoe question is $25 per canoe renatal.

OpenStudy (anonymous):

it was an accident

OpenStudy (anonymous):

thats what i got :)

OpenStudy (mathstudent55):

This? \( \large{8x^\frac{3}{7} } \)

OpenStudy (mathstudent55):

Is the 8 also raised to teh poer?

OpenStudy (anonymous):

its 8x 3/7

OpenStudy (mathstudent55):

Or this? \(\large{(8x)^\frac{3}{7} }\)

OpenStudy (anonymous):

8x^3/7

OpenStudy (mathstudent55):

If there are no parentheses, then only x is raised to the power; the 8 is not raised to the power.

OpenStudy (anonymous):

ya, no parenthesis

OpenStudy (mathstudent55):

\( \large{8x^\frac{3}{7} }\) \(\large{ = 8 \sqrt[7]{x^{3}} } \) When you have a fraction as a power, the numerator is an exponent and the denominator is the index of the root.

OpenStudy (anonymous):

ohh ok

OpenStudy (mathstudent55):

Answer is B.

OpenStudy (anonymous):

is the next question either a or b?

OpenStudy (mathstudent55):

What is the equation of the line that passes through the point (4, 1) and is perpendicular to the line y = –2x + 4 The slope of the given line is -2. That means the slope of the perpendicular line is 1/2. That eliminates A and B.

OpenStudy (anonymous):

oh i got -1/2 slope

OpenStudy (anonymous):

okk so c or d :3

OpenStudy (mathstudent55):

What are the slopes of perpendicular lines?

OpenStudy (anonymous):

umm?

OpenStudy (mathstudent55):

If you are given the slope of a line, what do you know about the slope of its perpendicular?

OpenStudy (anonymous):

is it half?

OpenStudy (mathstudent55):

No. The slopes of two perpendicular lines multiply to -1. The slopes of perpendicular lines are negative reciprocals. That means that if you know the slope of a line, to get the slope of its perpendicular, 1. write the slope of the line as a fraction, then 2. flip the fraction, then 3. change the sign.

OpenStudy (mathstudent55):

The given line is y = -2x + 4 What is the slope of this line?

OpenStudy (anonymous):

-2

OpenStudy (mathstudent55):

A line written in the slope-intercept form has the equation \( y = mx + b \) where m = slope and b = y-intercept

OpenStudy (anonymous):

ok :) -2 = slope

OpenStudy (mathstudent55):

Correct. m is the number that multiplies x in the equation of a line, and m is the slope, so in your problem, the slope is -2.

OpenStudy (mathstudent55):

Now to find the slope of the perpendicular, follow my 3 steps above. 1. write the slope as a fraction.

OpenStudy (anonymous):

-2/1

OpenStudy (mathstudent55):

Good. 2. flip the fraction

OpenStudy (anonymous):

1/-2

OpenStudy (mathstudent55):

Good. 3. change the sign.

OpenStudy (anonymous):

1/2

OpenStudy (mathstudent55):

Good. The slope we need is 1/2. We know the equation of a line is y = mx + b, where m = slope. Our slope is 1/2, so we can put 1/2 in form: \( y = \dfrac{1}{2}x + b \)

OpenStudy (anonymous):

y=4?

OpenStudy (anonymous):

i mean b=4

OpenStudy (mathstudent55):

Not necessarily. This is how we find b. We are given a point \( (x, y) = (4, 1) \) We input that point into x and y of the eqation and solve for b. \( 1 = \dfrac{1}{2}(4) + b \) \(1 = 2 + b\) \(-1 = b \) \(b = -1\) Now that we have b = -1, we can complete our equation: \( y = \dfrac{1}{2}x - 1 \) That is the equation we need. As you can see it's D.

OpenStudy (anonymous):

ohh ok :)

OpenStudy (anonymous):

btw i got a for the next one not b

OpenStudy (mathstudent55):

First I need to understand what this means. Solve 6the square root of x plus one) – 3 = 15

OpenStudy (mathstudent55):

\( \large{ 6 \sqrt{x + 1} - 3 = 15 } \) Is this the correct equation?

OpenStudy (anonymous):

6 sqrt x+1-3=15

OpenStudy (anonymous):

yes :)

OpenStudy (mathstudent55):

First, add 3 to both sides of the equation.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

is 8 right?

OpenStudy (mathstudent55):

\(\large{ 6 \sqrt{x + 1} - 3 = 15 } \) \(\large{ 6 \sqrt{x + 1} = 18 } \)

OpenStudy (mathstudent55):

Now divide both sides by 6, to get rid of the 6 multiplying the square root.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

\(\large{ \sqrt{x + 1} = 3 } \)

OpenStudy (mathstudent55):

Now you need to square both sides to get rid of the square root.

OpenStudy (mathstudent55):

\(\large{ (\sqrt{x + 1})^2 = (3)^2 } \) \(\large{ x + 1 = 9 } \) Subtract 1 from both sides: \(\large{ x = 8 } \)

OpenStudy (anonymous):

so i was right :D

OpenStudy (mathstudent55):

Now make sure 8 works in the original equation: \(\large{ 6 \sqrt{8 + 1} - 3 = 15 } \) \(\large{ 6 \sqrt{9} - 3 = 15 } \) \(\large{ 6 \times 3 - 3 = 15 } \) \(\large{ 18 - 3 = 15 } \) \(\large{ 15 = 15 } \) x = 8 is correct

OpenStudy (anonymous):

yay

OpenStudy (mathstudent55):

Answer is A

OpenStudy (anonymous):

ok, what about the next one?

OpenStudy (mathstudent55):

The last problem, you need to look at the graph. There are two inequalities graphed.

OpenStudy (anonymous):

is it b or c?

OpenStudy (anonymous):

i take that back

OpenStudy (anonymous):

4x-y is less than or equal to -10 ?

OpenStudy (mathstudent55):

One is a dashed line shaded blue under it. Form the y-intercept and slope its inequality is: \( y < \dfrac{1}{ 2}x +4 \) The other inequality has a solid line and is shaded orange underneath. From the y-intercept and slope the inequality is: y≤4x+10 Those are the two inequalities.

OpenStudy (anonymous):

c?

OpenStudy (mathstudent55):

y <= 4x + 10 -4x + y <= 10 4x - y >= -10

OpenStudy (mathstudent55):

You didn't show the choices with that question.

OpenStudy (anonymous):

2x – 5y greater than or equal to –20 4x – y > –10 2x – 5y > –20 4x – y greater than or equal to –10 2x – 5y < –20 4x – y less than or equal to –10 2x – 5y less than or equal to –20 4x – y < –10

OpenStudy (anonymous):

a?